OpenStudy (anonymous):

Write the given expression in terms of x and y only. sin(sin^−1 (x) + cos^−1 (y))

OpenStudy (anonymous):

sin(arcsin(x)) = x sin(arccos(x)) = sqrt(1-x^2) So... \[\sin(\arcsin(x) + \arccos(y)) = x + \sqrt{1-y^2}\] \[\sqrt{1-y^2} = -x \] \[y = plus or minus \sqrt{1-x^2}\]